Hi all, I am still having trouble with the following problem.
An object travels which velocity (3,1,2)
a) What is the speed of the object?
b) The object slows to speed 2 without changing direction. Give its velocity vector.
Thanks, -Nick.

Answer by Alan3354(69443)

An object travels which velocity (3,1,2)
a) What is the speed of the object?
Velocity is a vector that has speed and direction. It's a term that's often used when speed should be used.
The speed of the velocity vector (3,1,2) is the magnitude of the vector.
speed = sqrt(3^2 + 1^2 + 2^2)
speed = sqrt(14)
-------------------
b) The object slows to speed 2 without changing direction. Give its velocity vector.
This is the same direction but with magnitude (speed) of 2.
v2 = 2(3,1,2)/sqrt(14)
= (6/sqrt(14), 2/sqrt(14),4/sqrt(14))